Lou Barlow, a divisional manager for Sage Company, has an opportunity to manufacture and sell one of two new products for a five-year period. His annual pay raises are determined by his division’s return on investment (ROI), which has exceeded 23% each of the last three years. He has computed the cost and revenue estimates for each product as follows:

Product A Product B
Initial investment:
Cost of equipment (zero salvage value) $ 390,000 $ 585,000
Annual revenues and costs:
Sales revenues $ 420,000 $ 500,000
Variable expenses $ 185,000 $ 222,000
Depreciation expense $ 78,000 $ 117,000
Fixed out-of-pocket operating costs $ 90,000 $ 70,000

The company’s discount rate is 21%.

Click here to view Exhibit 7B-1 and Exhibit 7B-2, to determine the appropriate discount factor using tables.

Required:

1. Calculate the payback period for each product.

2. Calculate the net present value for each product.

3. Calculate the internal rate of return for each product.

4. Calculate the project profitability index for each product.

5. Calculate the simple rate of return for each product.

6a. For each measure, identify whether Product A or Product B is preferred.

6b. Based on the simple rate of return, Lou Barlow would likely:
Answer #1

1 payback period

= investment / annual cash flow

A B
ANNUAL REVENUE 420000 500000
VARIABLE EXPENSE (185000) (222000)
FIXEDOPERATING (90000) (70000)
ANNUAL CASH FLOW 145000 208000
LESS: DEPRECIATION (78000) (117000)
NET INCOME 67000 91000
PAYBACK 2.69[390000/145000] 2.81[585000/208000]

2.NET PRESENT VALUE

A B
ANNUAL CASH INFLOW 145000 208000
ANNUITY AT 21% 5 YEARS 2.926 2.926
PRESENT VALUE OF CASH FLOW 424270[145000*2.926] 608608[208000*2.926]
INITIAL OUTFLOW (390000) (585000)
NET PRESENT VALUE 34270$ 23608$

3.IRR WE WILL FIND OUT NPV AT TWO DIFERENT DISCOUNT RATE

AT 30%

ANNUAL CASH INFLOW 145000 208000
ANNUITY AT 25% 5 YEARS 2.6893 2.6893
PRESENT VALUE OF CASH FLOW 389949[145000*2.6893] 559374[208000*2.6893]
INITIAL OUTFLOW (390000) (585000)
NET PRESENT VALUE -51$ -25626

R1=21%

NPV 1=34270

NPV2 =-51

R2=25%

IRR = R1+NPV1(R2-R1)/(NPV1-NPV2)

=0.21 + 34270(0.25-0.21)/(34270- (-51)

=0.21+0.0399

24.99%

PROJECT A=25% AT 25% NPV WILL BE NEAR TO ZERO

PROJECY B

=0.21 + 23608(0.25-0.21)/(23608-(-25626)

=0.21+944.32/49234

=0.21+.0192

=23%

4.PROJECT PI

PRESENT VALUE OF CASH FLOW/ INVESTMENT

=424270/390000

PROJECT A=1.09

PROJECT B=608608/585000

PROJECT B=1.04

5.SIMPLE RATE OF RETURN

ANNUAL INCOME/ INVESTMENT

PROJECT A=67000/390000

=17.18%

B= 91000/585000

=15.56%

6a.

CHOOSE EXPLANATION
payback A A HAS LESSER PAYBACK PERIOD. INITIAL OUTFLOW WILL BE COVERED BACK SOONER BY A
NPV A BETTER NPV
IRR A HIGHER RATE OF RETURN
PROFITABILITY A HIGHER PI

6B.SIMPLE RATE OF RETURN A SHOULD BE SELECTED AS IT HAS HIGHER RATE OF RETURN.